Find the equation of the line that passes through the point (1,-2) and is perpendicular to the line passing through the points (-2,-3) and (4,3)
Answer by mananth(16946)


x1 y1 x2 y2
-2 -3 4 3

slope m = (y2-y1)/(x2-x1)
( 3 - -3 )/( 4 - -2 )
( 6 / 6 )
m= 1
slope of line perpendicular to the line joining these two points will be -1 ( m1*m2-1)
m= -1

Plug value of the slope and point (1, -2) in
Y = m x + b
-2 = -1 + b
b= -2 - -1
b= -1
So the equation will be
Y = - x -1
m.ananth@hotmail.ca